OpenStudy (dls):

A stone is dropped from the top of a cliff of height h . n seconds later a second stone is projected downwards from the same cliff with a vertical velocity u.The condition so that both the stones meet at the bottom of the cliff together is?

OpenStudy (ganpat):

x= Ut + 1/2 at2 .. U = initial velocity, a= Accelaration = g t = time required to reach ground.. Initial Energy is Potential energy, U = mgh.. As height and gravity remains same, Only factor changing will be Mass of the stone.. I guess, this can help u !!

OpenStudy (anonymous):

this is what i got so far: if stone 1 takes time t to reach the ground. then stone 2 should take time (t-n) for them both to meet at the bottom. so I got t = (gn^2 - 2un)/ (2u + 2gn) if we substitute this in h = -1/2 gt^2...........we should get a relation. but this seems to be complicated. anyone got better ideas?

OpenStudy (anonymous):

Let after n seconds first stone takes t seconds to reach the bottom. So first stone takes totally (n+t) seconds to reach bottom of the cliff . And then the second ston ]e takes totally t seconds to reach bottom (as both has to meet at the bottom) then the motion-eqn for first stone is: \[h=(1/2)g(t+n)^{2}...................(1)\](as it is simply dropped so no initial velocity,and g is acceleration due to gravity) and the motion-eqn for the second stone is: \[h=ut+(1/2)g.t^{2}...................(2)\](as it has initial velocity u) then adding (1) and (2) and separating h we get: \[(1/2)g.(t+n)^{2}=ut+(1/2)g.t^{2}\] \[(1/2)g[(t+n)^{2}-t^{2}]=ut\] \[(1/2)g[(2t+n).n]=ut\] \[2tn+n^{2}=2ut/g\] \[2t[(u/g)-n]=n^{2}\] \[t=(1/2)[n^{2}g/(u-ng)]\]

OpenStudy (anonymous):

@ganpat how can mass of the stone change? if we take that approach then we'll have to consider KE at the bottom. so mgh = 1/2 mv^2

OpenStudy (ganpat):

@Vaidehi09 : The Second stone can be of different mass.. This can be a parameter for them to meet ??

OpenStudy (anonymous):

@Vaidehi09 mass has nothing to do here in kinematics. I hope you know that all particle or object irrespective of their mass fall in same time from same height.

OpenStudy (anonymous):

but since mass is not mentioned in the prob. we can't have a relation in terms of it.

OpenStudy (anonymous):

@biswajit_paul yes, i'm aware.

OpenStudy (ganpat):

if not mass than, Area can be the factor... Surface are ??

OpenStudy (anonymous):

ok @Vaidehi09 so their is no problem if the masses of the stones are different. And obviously we have neglect air friction or other dssipative force .

OpenStudy (anonymous):

@Ganpat the surface area has nothing to do here as we have not considered friction.

OpenStudy (anonymous):

we need to find a relation in terms of h, u, n, g...most probably. coz these terms are specified in the question. so any extra terms we take while solving, need to be eliminated. so why would we add extra terms as mass and area here?

OpenStudy (anonymous):

hello..I have said you that mass is here useless term. Forget about it..

OpenStudy (anonymous):

i know. forget mass and area.

OpenStudy (anonymous):

umm...so what abt the answer here? @DLS what is it?

OpenStudy (anonymous):

@Vaidehi09 havent you seen my reply? I have given an answer with derivation. Whats about it?

OpenStudy (anonymous):

don't we have to eliminate t? since t is a variable chosen by us. it isn't given to us in the question.

OpenStudy (dls):

v(6n-9)/2n

OpenStudy (anonymous):

What I have found is also a condition..What do you think?

OpenStudy (dls):

@mukushla

OpenStudy (anonymous):

@DLS how did you find out the condition you have given here?

OpenStudy (dls):

That is the answer

OpenStudy (anonymous):

ok. let me try

OpenStudy (anonymous):

@DLS I hope that the v will be u. Isn't it?

OpenStudy (anonymous):

@biswajit_paul yes, that is a condition. I also got something along the same lines. but during substitution it got complicated.

OpenStudy (anonymous):

yes...that v should be u.

OpenStudy (anonymous):

be back in 15 mins!

OpenStudy (anonymous):

@DLS this is condition for what? h?

OpenStudy (anonymous):

hello @DLS

OpenStudy (anonymous):

@DLS no...that v should be u. or check the ques again. what is given in the ques - v or u? also...the answer...is it in the form of an equation or its just an expression?

OpenStudy (dls):

It is v only and answer is what i posted earlier only

OpenStudy (dls):

@mukushla @Vaidehi09 I'n not getting 2 in the denominator :/

OpenStudy (anonymous):

how did u proceed?

OpenStudy (dls):

i did from the beginning

OpenStudy (dls):

See,distance travelled in n seconds => sn = ut+1/2at^2 And distance travelled in nth second is given as: Sn = S(t-1) = u+a/2(2t-1) I usd this equation to get something good :o try it,maybe helps..

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

how come the answer is not in the form of an equation?

OpenStudy (anonymous):

i dont know if this is right or not \[n=\frac{v+\sqrt{2gh}-\sqrt{v^2+2gh}}{g}\]

OpenStudy (anonymous):

i got this eq for time t: \[t = gn ^{2} \div 2( v - gn )\]

OpenStudy (anonymous):

but the final ans has doesn't have g, t, or h.

OpenStudy (anonymous):

what is the final answer?

OpenStudy (anonymous):

v(6n - 9) / 2n what's more....its not even an equation!

OpenStudy (anonymous):

!!!!

OpenStudy (anonymous):

i know, right!

OpenStudy (dls):

Like I said,distance travelled in nth second is given as: Sn = S(t-1) = u+a/2(2t-1) use this and get the answer !!

OpenStudy (anonymous):

how did u get that eq?

OpenStudy (dls):

I got it in "Terms" of that :o im not substituting values right i guess :/

OpenStudy (anonymous):

u mean u derived it? how? show it here.

OpenStudy (dls):

I just have that as a formula :P

OpenStudy (anonymous):

i can't recognize it.....

OpenStudy (dls):

private formulas~

OpenStudy (dls):

arre just tell the time's value..will it be n-3?

OpenStudy (anonymous):

t is the time taken for stone1 to reach the ground ryt? so stone 2 will take t-n time.

OpenStudy (dls):

OH WTH SORRY THIS IS SOME OTHER QUES LOL SORRY :/

OpenStudy (anonymous):

hi friends @DLS could u plz check the final answer again ?

OpenStudy (dls):

ANSWER IS ! 8h(u-gn)^2 = gn^2(2u-gn)^2 sorry,i thought this was the next ques which i ws doing!

OpenStudy (anonymous):

@DLS ...................................................seriously, u had us racking our brains for sooooo long when we already had the answer.

OpenStudy (dls):

lol..sorry :/ i thought i posted the ques i was doing :/ im sorry..

OpenStudy (anonymous):

thats what i thought n=f(u,h)

OpenStudy (anonymous):

oh well....its fine already. so we solved this one too.

OpenStudy (dls):

arre solution :P ?

OpenStudy (anonymous):

ok...proceed this way: h = 0 + 0.5gt^2............stone 1..............=> (i) h = u(t-n) + 0.5g (t-n)^2............stone 2 then equate these 2....coz h is on LHS of both. get the equation for t substitute that in (i) u'll get the answer.

OpenStudy (anonymous):

hm, I have also got this answer.

OpenStudy (anonymous):

@biswajit_paul hi5!

OpenStudy (anonymous):

@Vaidehi09 yes